Two charges are located on the x axis: q1 = +6AfLC at x1 + 3.5 cm, and q2-+6.414c at x2 =-3.5 cm. Two other charges are located on the y axis: q3 = +5.0/c at y3 C at y4 = +7.1 cm. Find (a) the magnitude and (b) the direction of the net electric field at the origin. +4.6 cm, and q4 =-12μ 42 (a) Numb (b) The net electric field points the tolerance is +/-2%
Answer #1

As the charge particles q1 and q2 are same and also in the same axis so the net electric field due to these individual charges cancel out each other, i.e net electric field is only due to q3 and q4

E= electric field due to q4 - electric field due to q3

E = k ( q4/ r4^2 - q3/ r3^2)

E = 9*10^9* 10^-6* (12/ 0.071^2 - 5/0.046^2)

E = 1.58*10^5 N/C

Directed along +y axis

======

Do comment in case any doubt, will reply for sure.. Goodluck